# Etude: Piano Cover Generation with a Three-Stage Approach — Extract, strucTUralize, and DEcode

### v1

@@ Abstract
Piano cover generation aims to automatically transform a pop song into a piano arrangement. While numerous deep learning approaches have been proposed, existing models often fail to maintain structural consistency with the original song, likely due to the absence of beat-aware mechanisms or the difficulty of modeling complex rhythmic patterns. Rhythmic information is crucial, as it defines structural similarity (e.g., tempo, BPM) and directly impacts the overall quality of the generated music.
In this paper, we introduce Etude, a three-stage architecture consisting of Extract, strucTUralize, and DEcode stages. By pre-extracting rhythmic information and applying a novel, simplified REMI-based tokenization, our model produces covers that preserve proper song structure, enhance fluency and musical dynamics, and support highly controllable generation through style injection. Subjective evaluations with human listeners show that Etude substantially outperforms prior models, achieving a quality level comparable to that of human composers.
I Introduction
Piano cover generation—the artistic rearrangement of a song into a piano performance—is a popular form of musical expression that thrives on social media and has grown into a commercially viable field. Yet producing a high-quality piano cover remains challenging, requiring both analysis of the original song’s musical features (e.g., harmony and rhythm) and sufficient music theory knowledge to create a coherent, aesthetically pleasing arrangement.
@@ I Introduction
Figure 1: The architecture of our proposed Etude framework, which comprises three main components: an Extractor for the extract stage, a Beat-Detector for the structuralize stage, and a Decoder for the decode stage.
@@
Figure 2: The overall architecture of Etude, where an Extractor derives harmonic features, a pre-trained Beat-Detector provides the rhythmic framework for tokenization and de-tokenization, and a Decoder receives these inputs along with style prompts. The Extractor and Decoder are trained separately.
@@
In summary, by disentangling the learning of structure from the generation of notes, Etude produces structurally sound, stylistically controllable, and high-quality piano covers, demonstrating state-of-the-art performance in extensive evaluations where it substantially outperformed previous methods and achieved a quality closer to human-composed arrangements. All code and audio demonstrations are available on our project page. [^1] https://xiugapurin.github.io/Etude/
II Background
@@
III Methodology
Our proposed framework, Etude, is designed to address the key challenges in APCG, particularly in ensuring structural consistency and enabling stylistic control. The architecture, illustrated in Figure 2, modularizes the generation process into three distinct stages: Extract, Structuralize, and Decode.
III-A Extract Stage
The goal of the Extract stage is to overcome the information bottleneck found in prior two-stage models[2, 3], which often use simplified lead sheet[19] representations that lose significant musical nuance. To address this, we produce a rich and dense feature representation that captures all potentially salient events from the source audio. Our Extractor’s architecture is adapted from the AMT-APC[4], which is based on the hFT-Transformer[20]. We use the source audio as input but modify the model’s loss sampling parameter ($\theta^{\text{matrix}}$) to encourage the transcription of a dense map of musical events rather than a sparse, playable arrangement. The final output is a MIDI-like feature sequence that provides a comprehensive and detailed foundation for the Decoder.
III-B Structuralize Stage
The Structuralize stage is central to ensuring structural correctness. Running in parallel with the Extractor, it employs a pre-trained Beat-Transformer[21] to analyze the source audio and extract the precise timings of all beats and downbeats. This information forms a definitive rhythmic framework $F_{beat}$, containing tempo, time signature, and measure boundaries, that serves as an immutable structural ground truth for the entire process. This framework underpins our decoupled approach: it provides all necessary metrical information for the tokenization of symbolic data and the final reconstruction of a structurally coherent MIDI file from the model’s output (see Section III-D).
III-C Decode Stage
@@
  • Relative Polyphony: This measures the change in the average number of notes per unique time event. It reflects a shift in harmonic texture; for example, a high value indicates a transformation towards denser block chords, while a low value suggests a shift towards sparser textures like arpeggios.
  • Relative Rhythmic Intensity: This measures the change in the density of rhythmic events over time. A higher value results in a more complex and rhythmically active phrase.
  • Relative Note Sustain: This measures the change in the average note duration, controlling the output’s articulation character from legato (high) to staccato (low).
These relative attributes are calculated for every bar pair in the training set and discretized into a small number of bins based on their statistical distribution. During training and inference, each token in a bar-pair sequence is assigned the three corresponding style bin IDs for that bar. These IDs are then passed through separate embedding layers, concatenated, and projected by a linear layer to form a single style embedding vector. This style embedding is added to the token and class ID embeddings to form the final input to the Transformer Decoder. This mechanism effectively conditions the entire generation process on the desired stylistic transformation, allowing for fine-grained, bar-level control over the final output’s characteristics.
Figure 3: An example of Tiny-REMI tokenization, where the sequence is delimited by Bar [BOS] and Bar [EOS] tokens, and each note event group (Note + Dur) is preceded by a Pos token marking its relative position.
III-D Data Representation
@@
IV Experiments
TABLE I: Comprehensive evaluation results of human composers, our proposed Etude models, and other baselines on both objective and subjective metrics. For objective metrics, we also show the delta ($\Delta$) from the human performance. For subjective metrics, higher is better ($\uparrow$).
@@ IV-C Baselines
  • Etude - Default: The Decoder’s output using default, median style attributes.
  • Etude - Prompted: The Decoder’s output with manually selected style prompts for each song. [^3] While our model supports per-measure style injection, for this evaluation, a single set of style vectors was applied to the entire piece.
IV-D Objective Metrics
@@
  • Warp Path Deviation (WPD) evaluates the structural similarity between the generated cover and the original song. This metric improves upon the WP-std from Music2MIDI[5] by calculating the standard deviation of the residuals from a linear regression of the DTW warp path. This approach makes WPD robust to global tempo variations while remaining sensitive to significant structural misalignments.
  • Rhythmic Grid Coherence (RGC) is introduced to evaluate rhythmic fluency and detect timing jitter. It quantifies how well the inter-onset intervals (IOIs) of a performance adhere to a consistent underlying metrical grid by first identifying a base tempo unit ($\tau$) from the most frequent IOIs, and then calculating the average deviation from the grid defined by $\tau$. A lower RGC score indicates a more rhythmically precise and coherent performance.
  • IOI Pattern Entropy (IPE) assesses the dynamic expressiveness and complexity of rhythmic patterns. This metric calculates the Shannon entropy of the distribution of overlapping n-grams within the generated performance’s IOI sequence, after the IOIs have been converted into a series of discrete symbols via k-means clustering. A higher IPE score suggests more varied and less repetitive rhythmic motifs.
IV-E Subjective Evaluation
To complement our objective metrics, we conducted a subjective listening test with 101 volunteers, who were categorized into three groups based on their musical experience: amateurs (<<3 years), intermediate (3-10 years), and expert (>>10 years). Each participant was asked to listen to the original version of a song, followed by seven piano cover versions presented in a randomized and anonymous order. These versions included the human performance and the outputs from the six models listed in Section IV-C. To ensure fairness, all model-generated MIDI files and the human performance (which was first transcribed to MIDI) were synthesized into audio using the same piano synthesizer. All audio excerpts were 60-second clips extracted from the beginning of each song.
@@ IV-E Subjective Evaluation
IV-F Results and Discussion
IV-F1 Evaluation Results
The comprehensive results for both objective and subjective evaluations are presented in Table I.
@@ IV-F1 Evaluation Results
The subjective evaluation results in Table I confirm the superiority of our proposed framework. Statistical analysis reveals that all three Etude variants significantly outperform all three baseline models (PiCoGen2, AMT-APC, and Music2MIDI) in Overall Quality (OVL), with all planned pairwise comparisons yielding a corrected $p<0.001$.
Diving into the details, the Human performance unsurprisingly achieved the highest scores across all criteria, establishing a clear upper bound. Among the AI models, both versions of our Etude Decoder were the top performers, with Etude - Default achieving the highest OVL score of 3.50. Notably, our decoder versions also lead in Fluency (FL) and Dynamic Expression (DE), indicating that their outputs are perceived as more natural and musical. Meanwhile, the Etude Extractor achieved the highest Similarity (SI) score among all models, confirming its effectiveness at capturing a dense and accurate representation of the source material. These results provide strong evidence that the Etude framework represents a substantial advancement in the quality of automatic piano cover generation.
TABLE II: Descriptive statistics of the Overall (OVL) scores for participant groups based on musical experience. Scores are presented as mean $\pm$ standard deviation.
@@
IV-F2 Analysis of Rater Experience
To validate our subjective findings, we analyzed the OVL scores by rater experience (Table II). While a trend emerged where more experienced listeners gave harsher scores, a statistical test showed that the difference among the groups was not significant ($p=0.070$). This suggests our main results are robust across listeners with varying musical expertise.
IV-F3 Ablation Study
@@
To isolate the impact of the model architecture from the influence of the dataset, we first trained a new version of the PiCoGen2 model, hereafter referred to as PiCoGen2*, following its official implementation but using our own high-quality dataset (detailed in Section IV-A). We then conducted a separate subjective listening test with 30 new participants. The experimental setup was identical to our main evaluation (Section IV-E), including the audio synthesis process and the four criteria (SI, FL, DE, OVL). The test compared the following three systems:
@@
The results of this ablation study are presented in Table III. The analysis reveals two key findings. First, while retraining the baseline on our dataset (PiCoGen2*) shows a slight improvement over the original PiCoGen2, the difference in OVL scores was not statistically significant ($p=0.0656$). Second, and more crucially, our Etude - Default model significantly outperformed PiCoGen2* ($p=0.0007$), even when both were trained on the exact same dataset. This provides strong evidence that while our data processing pipeline contributes positively, the primary driver for the substantial performance gain is the architectural innovations of the Etude framework itself.
TABLE III: Subjective evaluation results for the ablation study, comparing the effects of the dataset and model architecture.
@@ V Conclusion
For future work, we observe that the performance of our framework is fundamentally upper-bounded by the capabilities of its front-end components: the Beat-Detector and the Extractor. The framework’s structural accuracy is limited by the precision of the beat tracker, while the Extractor’s process of flattening the source audio into a single feature stream can lead to information loss. This limitation is evident in our subjective results (Table I), where both Etude Decoder versions scored markedly lower on Similarity (SI) than the direct Etude Extractor output. We hypothesize this is because the flattened features make it difficult for the model to identify the primary melody of the original song, causing the Decoder to generate incomplete melodic lines. Future work could thus explore both the integration of more advanced beat-tracking modules and the development of multi-stream extractors capable of disentangling different musical sources. Finally, our work highlights the persistent gap between current objective metrics and human perception. Developing learnable, perception-aligned evaluation models could provide a more reliable and automated benchmark for the field. Overall, the Etude framework provides a robust foundation for generating high-quality, controllable piano covers, paving the way for more musical and interactive AI-powered creative tools.

### v2

@@ Abstract
Piano cover generation aims to automatically transform a pop song into a piano arrangement. While numerous deep learning approaches have been proposed, existing models often fail to maintain structural consistency with the original song, likely due to the absence of beat-aware mechanisms or the difficulty of modeling complex rhythmic patterns. Rhythmic information is crucial, as it defines structural similarity (e.g., tempo, BPM) and directly impacts the overall quality of the generated music. In this paper, we introduce Etude, a three-stage architecture consisting of Extract, strucTUralize, and DEcode stages. By pre-extracting rhythmic information and applying a novel, simplified REMI-based tokenization, our model produces covers that preserve proper song structure, enhance fluency and musical dynamics, and support highly controllable generation through style injection. Subjective evaluations with human listeners show that Etude substantially outperforms prior models, achieving a quality level comparable to that of human composers.
Index Terms—  Automatic Piano Cover Generation, Music Generation, Music Information Retrieval, Automatic Music Transcription
1 Introduction
Piano cover generation—the artistic rearrangement of a song into a piano performance—is a popular form of musical expression that thrives on social media and has grown into a commercially viable field, with applications ranging from enabling learners to obtain customized sheet music to assisting creators in producing arrangements for streaming or distribution. Yet producing a high-quality piano cover remains challenging, requiring both analysis of the original song’s musical features (e.g., harmony and rhythm) and sufficient music theory knowledge to create a coherent, aesthetically pleasing arrangement.
@@ 1 Introduction
Fig. 1: The architecture of our proposed Etude framework, which comprises three main components: an Extractor for the extract stage, a Beat-Detector for the structuralize stage, and a Decoder for the decode stage.
@@
Fig. 2: The overall architecture of Etude, where an Extractor derives harmonic features, a pre-trained Beat-Detector provides the rhythmic framework for tokenization and de-tokenization, and a Decoder receives these inputs along with style prompts. During training, the lower piano cover branch provides ground-truth target sequences ($Y$) aligned via weak-alignment. During inference, the lower branch content is replaced by autoregressively generated outputs: the Decoder generates each bar $Y_{i}$ from the corresponding feature bar $X_{i}$ and previous context. The Extractor and Decoder are trained separately.
@@
In summary, by disentangling the learning of structure from the generation of notes, Etude produces structurally sound, stylistically controllable, and high-quality piano covers, demonstrating state-of-the-art performance in extensive evaluations where it substantially outperformed previous methods and achieved a quality closer to human-composed arrangements. All code and audio demonstrations, along with a full version of the manuscript, are available on our project page. [^1] https://xiugapurin.github.io/Etude/
2 Background
@@
3 Methodology
Our proposed framework, Etude, is designed to address the key challenges in APCG, particularly in ensuring structural consistency and enabling stylistic control. The architecture, illustrated in Figure 2, modularizes the generation process into three distinct stages: Extract, Structuralize, and Decode. This section details the motivation and implementation of each stage.
3.1 Extract Stage
The goal of the Extract stage is to overcome the information bottleneck found in prior two-stage models[19, 18], which often use simplified lead sheet[5] representations that lose significant musical nuance. To address this, we produce a rich and dense feature representation that captures all potentially salient events from the source audio. Our Extractor’s architecture is adapted from AMT-APC[12], which is a piano cover generation model fine-tuned from the hFT-Transformer[21] for automatic music transcription. We use the source audio as input but modify the model’s loss sampling parameter ($\theta^{\text{matrix}}$) to encourage the transcription of a dense map of musical events rather than a sparse, playable arrangement. The final output is a MIDI-like feature sequence that provides a comprehensive and detailed foundation for the Decoder.
3.2 Structuralize Stage
The Structuralize stage is central to ensuring structural correctness. Running in parallel with the Extractor, it employs a pre-trained Beat-Transformer[26] to analyze the source audio and extract the precise timings of all beats and downbeats. This information forms a definitive rhythmic framework $F_{beat}$, containing tempo, time signature, and measure boundaries, that serves as an immutable structural ground truth for the entire process. This framework underpins our decoupled approach: it provides all necessary metrical information for the tokenization of symbolic data and the final reconstruction of a structurally coherent MIDI file from the model’s output (see Section 3.4 for details on how $F_{beat}$ is used in tokenization and de-tokenization).
3.3 Decode Stage
@@
During inference, the bar-wise mixing structure is maintained, but the target sequences Y come from the model’s own generations rather than ground-truth piano covers. Specifically, the generation proceeds bar-by-bar: for each bar $i$, the Decoder receives the current feature bar $X_{i}$ along with up to four preceding $(X,Y)$ pairs as context. The model then autoregressively generates tokens for $Y_{i}$ until a bar-end token is produced. The completed pair $(X_{i},Y_{i})$ is appended to the context window for generating the next bar. This approach ensures that the model leverages its own outputs to maintain long-term musical coherence during generation.
@@
  • Relative Polyphony: change in the average number of notes per unique time event. It reflects a shift in harmonic texture; for example, a high value indicates a transformation towards denser block chords, while a low value suggests a shift towards sparser textures like arpeggios.
  • Relative Rhythmic Intensity: change in the density of rhythmic events over time. A higher value results in a more complex and rhythmically active phrase.
  • Relative Note Sustain: change in the average note duration, controlling the output’s articulation character from legato (high) to staccato (low).
These relative attributes are calculated for every bar pair in the training set and discretized into three bins 0, 1, 2 (low, medium, high) based on their statistical distribution. During training and inference, each token in a bar-pair sequence is assigned the three corresponding style bin IDs for that bar. These IDs are then passed through separate embedding layers, concatenated, and projected by a linear layer to form a single style embedding vector. This style embedding is added to the token and class ID embeddings to form the final input to the Transformer Decoder. This mechanism effectively conditions the entire generation process on the desired stylistic transformation, allowing for fine-grained, bar-level control over the final output’s characteristics.
Fig. 3: An example of Tiny-REMI tokenization, where the sequence is delimited by Bar [BOS] and Bar [EOS] tokens, and each note event group (Note + Dur) is preceded by a Pos token marking its relative position.
3.4 Data Representation
@@
4 Experiments
Table 1: Comprehensive evaluation results of human composers, our proposed Etude models, and other baselines on both objective and subjective metrics. For objective metrics, we also show the delta ($\Delta$) from the human performance. For subjective metrics, higher is better ($\uparrow$).
@@ 4.3 Baselines
  • Etude - Default: The Decoder’s output using the default (median) style attributes: (1, 1, 1) for polyphony, rhythmic intensity, and note sustain.
  • Etude - Prompted: The Decoder’s output with manually selected style prompts for each song. [^3] The selected style attributes are based on the author’s preference. While our model supports per-measure style injection, for this evaluation, a single set of style vectors was applied to the entire piece.
4.4 Objective Metrics
@@
  • Warp Path Deviation (WPD) evaluates structural similarity between the generated cover and the original song. Given a DTW warp path $(S_{1},P_{1}),...,(S_{n},P_{n})$} obtained from chroma features via MrMsDTW[17], we fit a linear regression $\hat{S_{k}}=a\cdot P_{k}+b$ to capture the global tempo trend, then compute WPD as the standard deviation of residuals:
    $\mathrm{WPD}=\sqrt{\frac{1}{L}\sum_{k=1}^{L}\left(S_{k}-\hat{S_{k}}\right)^{2}}$
    This approach improves upon WP-std from Music2MIDI[25] by remaining robust to global tempo variations while detecting structural misalignments such as missing bars or misinterpreted chords.
  • Rhythmic Grid Coherence (RGC) evaluates rhythmic fluency by quantifying how well inter-onset intervals (IOIs) adhere to a metrical grid. We extract all unique onset times to compute the IOI sequence, identify the most frequent IOIs as the core rhythm set $I$, then search for the optimal base tempo unit $\tau\in I$ that minimizes grid deviation:
    $\mathrm{RGC}=\min_{\tau\in\mathcal{I}}\left(\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\left|\frac{i}{\tau}-\text{round}\left(\frac{i}{\tau}\right)\right|\right)$
    A lower RGC indicates more precise alignment to a consistent metrical grid, reflecting higher rhythmic coherence.
  • IOI Pattern Entropy (IPE) assesses the dynamic complexity of rhythmic patterns. We extract the IOI sequence, apply log-transformation and k-means clustering to discretize IOIs into symbols, then compute the Shannon entropy over all overlapping n-grams:
    $\mathrm{IPE}=-\sum_{g\in G_{n}}P(g)\log_{2}P(g)$
    where $G_{n}$ is the set of unique n-gram patterns. A higher IPE suggests more varied and less repetitive rhythmic motifs, balancing between monotonous (low IPE) and chaotic (excessively high IPE) patterns.
4.5 Subjective Evaluation
To complement our objective metrics, we conducted a subjective listening test with 101 volunteers, who were categorized into three groups based on their musical experience: amateurs (<<3 years), intermediate (3-10 years), and expert (>>10 years). Each participant was asked to listen to four randomly selected songs. For each song, they first heard the original version, followed by seven piano cover versions presented in a randomized and anonymous order. These versions included the human performance and the outputs from the six models listed in Section 4.3. To ensure fairness, all model-generated MIDI files and the human performance (which was first transcribed to MIDI) were synthesized into audio using the same piano synthesizer. All audio excerpts were 60-second clips extracted from the beginning of each song.
@@ 4.5 Subjective Evaluation
4.6 Results and Discussion
4.6.1 Evaluation Results
The comprehensive results for both objective and subjective evaluations are presented in Table 1.
@@ 4.6.1 Evaluation Results
The subjective evaluation results in Table 1 confirm the superiority of our proposed framework. Statistical analysis reveals that all three Etude variants significantly outperform all three baseline models (PiCoGen2, AMT-APC, and Music2MIDI) in Overall Quality (OVL), with all planned pairwise comparisons yielding a corrected $p<0.001$.
Diving into the details, the Human performance unsurprisingly achieved the highest scores across all criteria, establishing a clear upper bound. Among the AI models, both versions of our Etude Decoder were the top performers, with Etude - Default achieving the highest OVL score of 3.50. Notably, our decoder versions also lead in Fluency (FL) and Dynamic Expression (DE), indicating that their outputs are perceived as more natural and musical. Interestingly, Etude - Default slightly outperformed Etude - Prompted. This is likely because, for fairness in evaluation, we applied fixed style vectors across all bars in the Prompted setting, which may have limited dynamic expressivity compared to the Default’s median attributes. Meanwhile, the Etude Extractor achieved the highest Similarity (SI) score among all models, confirming its effectiveness at capturing a dense and accurate representation of the source material. These results provide strong evidence that the Etude framework represents a substantial advancement in the quality of automatic piano cover generation.
Table 2: Descriptive statistics of the Overall (OVL) scores for participant groups based on musical experience. Scores are presented as mean $\pm$ standard deviation.
@@
4.6.2 Analysis of Rater Experience
To validate our subjective findings, we analyzed the OVL scores by rater experience (Table 2). While a trend emerged where more experienced listeners gave harsher scores, a statistical test showed that the difference among the groups was not significant ($p=0.070$). This suggests our main results are robust across listeners with varying musical expertise.
4.6.3 Ablation Study
@@
To isolate the impact of the model architecture from the influence of the dataset, we first trained a new version of the PiCoGen2 model, hereafter referred to as PiCoGen2*, following its official implementation but using our own high-quality dataset (detailed in Section 4.1). We then conducted a separate subjective listening test with 30 new participants. The experimental setup was identical to our main evaluation (Section 4.5), including the audio synthesis process and the four criteria (SI, FL, DE, OVL). The test compared the following three systems:
@@
The results of this ablation study are presented in Table 3. The analysis reveals two key findings. First, while retraining the baseline on our dataset (PiCoGen2*) shows a slight improvement over the original PiCoGen2, the difference in OVL scores was not statistically significant ($p=0.0656$). Second, and more crucially, our Etude - Default model significantly outperformed PiCoGen2* ($p=0.0007$), even when both were trained on the exact same dataset. This provides strong evidence that while our data processing pipeline contributes positively, the primary driver for the substantial performance gain is the architectural innovations of the Etude framework itself.
Table 3: Subjective evaluation results for the ablation study, comparing the effects of the dataset and model architecture.
@@ 5 Conclusion
For future work, we observe that the performance of our framework is fundamentally upper-bounded by the capabilities of its front-end components: the Beat-Detector and the Extractor. The framework’s structural accuracy is limited by the precision of the beat tracker, while the Extractor’s process of flattening the source audio into a single feature stream can lead to information loss. This limitation is evident in our subjective results (Table 1), where both Etude Decoder versions scored markedly lower on Similarity (SI) than the direct Etude Extractor output. We hypothesize this is because the flattened features make it difficult for the model to identify the primary melody of the original song, causing the Decoder to generate incomplete melodic lines. Future work could thus explore both the integration of more advanced beat-tracking modules and the development of multi-stream extractors capable of disentangling different musical sources. Finally, our work highlights the persistent gap between current objective metrics and human perception. Developing learnable, perception-aligned evaluation models could provide a more reliable and automated benchmark for the field. Overall, the Etude framework provides a robust foundation for generating high-quality, controllable piano covers, paving the way for more musical and interactive AI-powered creative tools.